OpenStudy (anonymous):

Solve x^2 + 8x – 48 = 0 by completing the square. Show your work for full credit. Solve x^2 + 8x – 48 = 0 by completing the square. Show your work for full credit. @Mathematics

OpenStudy (anonymous):

(x+8/2)^2-(8/2)^2-48=0 (x+4)^2-64=0 (x+4)^2=64 x+4=8 or -8 x=12 or -4

OpenStudy (anonymous):

Thank you (: